OpenStudy (anonymous):

Use Linear approximation to estimate the following quantities. Choose a value of a to produce a small errror. \[\cos31 \] Linear approx form \[\ L(x) = f(a)+f \prime(a)(x-a) \]

OpenStudy (anonymous):

My issue is figuring out how to get the answer using trig functions. According to my book, as I am evaluating my values using the linear approx. formula, my answers have to be in radians. So I want to choose 30 as my a, \[\cos\frac{ pi }{ 6 }=.866\] ^ This is my value that approximating Now plugging in pi/6 in the formula \[\cos\frac{ pi }{ 6 }+(-sin\frac{ pi }{ 6 })(31-30)\]I get .366. Where is my mistake?

OpenStudy (phi):

I don't think you should mix in degrees and radians. in other words, change 31-30 = 1deg into radians = \( \frac{\pi}{180} \)

OpenStudy (anonymous):

Ah, I see now i got the right answer. Thanks. So just to make sure i'm doing this right, if i have say tan3, and an a value of 2... \[\tan3 = tanpi/60=.05\] \[\tan(\frac{\pi}{90})+\sec^2(\frac{\pi}{90})(\frac{\pi}{180})=.035=.04\]

OpenStudy (phi):

if you have tan 3º as your starting point , your formula should be \[ f(x) \approx \tan 3º + sec^2( 3º )(x - 3) \cdot \frac{\pi}{180}\] so you should get minus pi/180 (not plus pi/180) in other words \[ \tan(\frac{\pi}{90})-\sec^2(\frac{\pi}{90})(\frac{\pi}{180}) \]

OpenStudy (anonymous):

my a is , 3 is the x value

OpenStudy (anonymous):

*a is 2.

OpenStudy (phi):

if a is 2 then you want tan(x) = tan(2º) + sec^2(2º) (x - 2º) * pi/180

OpenStudy (phi):

so your equation looks ok (but I would use more decimals)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thanks.

OpenStudy (phi):

tan(3º) = 0.052407779 the approximation is \[ \tan(3º) \approx 0.034920769 + 1.00121946 \cdot \frac{\pi}{180} \\ = 0.034920769 + 1.00121946 \cdot 0.017453292 \\= 0.052395 \] which is close to the actual value 0.052407779